a ballstrikes a rough plane with velocity v and rebounds with velocity v/2. if angle of reflection is equal to angle of incidence then coefficient of friction is???


ans. tan@/3

Grade:

1 Answers

Chetan Mandayam Nayakar
13 years ago

vertical impulse=(3v/2)cos@, horizontal impulse=u(3v/2)cos@=(v/2)sin@ implies u=tan@/3